A family has two cars. The first car has a fuel efficiency of 35 miles per gallon of gas and the second has a fuel efficiency of 40 miles per gallon of gas. During one particular week, the two cars went a combined total of 1850 miles, for a total gas consumption of 50 gallons. How many gallons were consumed by each of the two cars that week?
Answer by Boreal(15235)

x gallons used by first car
50-x gallons used by the second car
35x+(50-x)40=1850, mileage equation
35x+2000-40x=1850
-5x=-150
x=30 gallons used by first car or 1050 miles
50-x=20 gallons used by second car or 800 miles
They add to 1850 miles